OpenStudy (anonymous):

Find the nonpermissible replacement for the variable in this expression. 5x/6x+3

OpenStudy (anonymous):

5x/6x+3 = undefined 1/6x+3 = undefined 6x+3 = 0 6x = -3 x = -1/2 x = -0.5

OpenStudy (anonymous):

thanks agian

OpenStudy (anonymous):

u r welcome :)